OpenStudy (anonymous):

Find the exact real value of arccos((square root of 2)/2)

OpenStudy (anonymous):

\[\arccos \left( \frac{ \sqrt{2} }{2 } \right)\]

OpenStudy (haseeb96):

arccos x = cos^-1 (x) now what you will do . just use your calculator first press shift button then go to the cos^-1 sign button then press these values you willl get your answer

OpenStudy (anonymous):

But how do I do it without the use of calculator? :)

OpenStudy (haseeb96):

well first you will divide sqrt 2 and 2 you will get 0.706 and now as you know the table of trigonometry there is cos 45= 0.707 So, we get it will be at 45 degrees.

OpenStudy (anonymous):

45 degrees or pi/4, right?

OpenStudy (haseeb96):

45 is degree, but when we express its answer in radians we will say pi/4 Good job girl @mathway

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Thank you @Haseeb96 !

OpenStudy (haseeb96):

Most welcome @mathway I am glad that i helped you in your problem :)

OpenStudy (haseeb96):

Do you have more questions ? if so tag me

OpenStudy (anonymous):

I'll open a new question. :)

OpenStudy (haseeb96):

fine